Jarett & Sons's common stock currently trades at $37.00 a share. It is expected to pay an annual dividend of $2.25 a share at the end of the year (D1 = $2.25), and the constant growth rate is 4% a year.

  1. What is the company's cost of common equity if all of its equity comes from retained earnings? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
  2. If the company issued new stock, it would incur a 19% flotation cost. What would be the cost of equity from new stock? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
Answer #1

a.Cost of equity=(D1/Current price)+Growth rate

=(2.25/37)+0.04

=10.08%(Approx).

b.Cost of equity=(D1/Current price(1-Flotation cost)+Growth rate

=(2.25/37(1-0.19))+0.04

=(2.25/29.97)+0.04

=11.51%(Approx).